### v1

# Cell divisions imprint long lasting elastic strain fields in epithelial tissues

Abstract
A hallmark of biological tissues, viewed as complex cellular materials, is the active generation of mechanical stresses by cellular processes, such as cell divisions. Each cellular event generates a force dipole that deforms the surrounding tissue. Therefore, a quantitative description of these force dipoles, and their consequences on tissue mechanics, is one of the central problems in understanding the overall tissue mechanics. In this work we analyze previously published experimental data on fruit fly D. melanogaster wing epithelia to quantitatively describe the deformation fields induced by a cell-scale force dipole. We find that the measured deformation field can be explained by a simple model of fly epithelium as a linearly elastic sheet. This fact allows us to use measurements of the strain field around cellular events, such as cell divisions, to infer the magnitude and dynamics of the mechanical forces they generate. In particular, we find that cell divisions exert a transient isotropic force dipole field, corresponding to the temporary localisation of the cell nucleus to the tissue surface during the division, and traceless-symmetric force dipole field that remains detectable from the tissue strain field for up to about $3.5$ hours after the division. This is the timescale on which elastic strains are erased by other mechanical processes and therefore it corresponds to the tissue fluidization timescale. In summary, we have developed a method to infer force dipoles induced by cell divisions, by observing the strain field in the surrounding tissues. Using this method we quantitatively characterize mechanical forces generated during a cell division, and their effects on the tissue mechanics.
I Introduction
How mechanical forces influence biological tissues is one of the central problems in animal development and regeneration Heisenberg and Bellaïche (2013); Brugués et al. (2014); LeGoff and Lecuit (2016); Jülicher and Eaton (2017); Tetley et al. (2019). Biological tissues are often described as soft viscoelastic materials Aigouy et al. (2010) that are elastic on sufficiently short timescales. Tissue fluidization on longer timescales arises from cell divisions, extrusions and intercalations, that restructure the tissue over time Ranft et al. (2010); Etournay et al. (2015); Curran et al. (2017); Kim et al. (2021).
Cell divisions are an intrinsic source of mechanical force generation within tissues. When a cell divides, particularly in a closely packed epithelium, it induces deformation in the surrounding tissue Gibson et al. (2011); Mao et al. (2013), potentially leading to cellular rearrangements Firmino et al. (2016) and contributing to the overall tissue flow Ranft et al. (2010); Matoz-Fernandez et al. (2017); Bocanegra-Moreno et al. (2023). Furthermore, cell divisions have been indicated to control the glassy cell dynamics observed in cultured epithelial tissues Angelini et al. (2011); Park et al. (2015) and explanted embryonic tissues Schotz et al. (2013). However, despite the fact that cell divisions are a hallmark of biological matter and seem to be major factor in controlling material properties of biological tissues, quantitative understanding of the mechanical forces generated by cell divisions and their immediate consequences on the surrounding tissue is still lacking.
In order to study the mechanics of cell divisions, we need to probe mechanical stresses in the surrounding tissue. One of the most common methods of probing mechanical stresses is the laser ablation experiment, in which one or multiple cells are destroyed by a laser. The resulting displacement of the ablated tissue boundary reflects the pre-stresses existing in the tissue before the ablation Farhadifar et al. (2007); Mayer et al. (2010); Bonnet et al. (2012); Kasza et al. (2014); Etournay et al. (2015); Piscitello-Gómez et al. (2023). Information about the pre-stresses are extracted either from the initial dynamics of tissue boundary displacement Mayer et al. (2010); Kasza et al. (2014) or from the final relaxed shape of the tissue Farhadifar et al. (2007); Bonnet et al. (2012); Dye et al. (2021); Piscitello-Gómez et al. (2023). Previous analyses focused on changes in position or shape of specific points or outlines in the vicinity of the ablation. Therefore, they could not provide an insight into the strain field induced in the surrounding tissue, and they had to rely on assumptions about tissue mechanical properties. Furthermore, to extract information about all components of the pre-stress in the tissue, previous methods require a precise, circular laser ablation Bonnet et al. (2012); Dye et al. (2021). If we could instead infer force dipoles induced by cellular processes such as cell divisions, we would be able to extract information about tissue mechanics in a completely non-invasive way. Such a method was recently developed to infer the properties of the core of plastic particle rearrangements in computer glasses Moriel et al. (2020, 2024). Therefore, the starting point of our work, on the way to study cell division mechanics, is to develop a method that allows us infer the force dipole in an epithelial tissue from the tissue strain field.
In this work, we first provide a detailed quantitative analysis of the tissue displacement field following a linear laser ablation performed in the pupal wing epithelium of the fruit fly D. melanogaster. We show that the response of the fly wing epithelium to a laser ablation is consistent with that of a two-dimensional linear elastic sheet. This allows us to infer the force dipole induced by the ablation, normalized by the elastic constant, from the observed strain field. Furthermore, by analyzing the dynamics of the strain field, we show that the dominant mode of dissipation in the fly wing epithelium on short timescales of $1s$ to $50s$ is the viscous dissipation, as opposed to frictional dissipation with a substrate.
Using our method to infer force dipole from its strain field, we determine the force dipole generated by cell divisions in the D. melanogaster wing disc. We find that the isotropic component of the cell division force dipole tensor is only transient and vanishes almost immediately after the division, indicating that no net cell growth is accumulated during the cell division. Furthermore, we find that the traceless-symmetric component of the force dipole, imprinted by the cell division, remains visible in the surrounding tissue for up to about $3.5$ hours after the division. This provides a measurement of the fluidization timescale in the tissue, and demonstrates importance of elastic interactions in tissue-scale morphogenetic processes that take place on a comparable timescale, such as wing disc eversion Fuhrmann et al. (2023).
II The fruit fly wing epithelium is a 2D linear elastic sheet
II.1 Linear laser ablation experiments
Linear laser ablation experiments, during which a small linear segment of tissue is destroyed by a laser, have been performed in the D. melanogaster wing epithelium during pupal morphogenesis Piscitello-Gómez et al. (2023). To obtain information about the pre-stress in the tissue, the dynamics of the closest cell-cell junction was measured as it retracted away from the ablation. Here, we use movies of laser ablation experiments published in Ref. Piscitello-Gómez et al. (2023) to study the deformation in the tissue surrounding the laser ablation. An example of such a laser ablation experiment is shown in Fig. 1 (A.i), where the red line indicates the line ablated by the laser. Following the ablation, the tissue relaxes until reaching a mechanical equilibrium after at most $T_{e}=50s$ in Fig. 1 (A.ii), see Methods.
To measure the tissue strain field, we segmented the apical cell surfaces over time and quantified the area $a$ and the elongation tensor $Q_{ij}$ of each cell, which quantifies its shape Etournay et al. (2015); Merkel et al. (2017). Since no cell divisions or rearrangements appear in the region of interest following the laser ablation, we quantify the tissue strain field ${U}_{ij}(\bm{r},t)$ from the change in cell shape and size relative to their values before the ablation, see Methods. The strain field tensor consists of the isotropic strain component $U_{kk}$ and traceless-symmetric component $\tilde{U}_{ij}$ which we denote as pure shear strain component. In Fig. 1 (A.iv) we show the pure shear strain field component $\tilde{U}_{xy}$ at $100s$ after the laser ablation, which exhibits a clear spatial pattern, see SI Fig. 2 (A) for the other traceless-symmetric component and isotropic component of experiment 1 and SI Fig. 3 (A) for all components of the experiment 2.
Our goal is to find the relationships between the induced strain field and the forces induced by the ablation. Since the process of ablation does not introduce any net force, the dominant component of the induced force field will be dipolar. Therefore, we aim to relate components of the strain field at distance $\vec{r}$ from the ablation to the components of the force dipole tensor, as illustrated in Fig. 1 (B.i,ii).
We first consider tissue strain after it has reached the mechanical equilibrium after the time $T_{e}$ and calculate the discrete angular Fourier spectrum $\tilde{U}^{ss}_{ij}(r,m)$, where and $r=|\bm{r}|$ is distance from the center of ablation and $m$ is the angular mode index, see Methods. We find the strongest signal in the second and the fourth mode of the pure shear strain components, see Fig. 1 (C.i) and SI Fig. 2 (B.ii), which correspond to $\cos(2\varphi)$ and $\cos(4\varphi)$ angular Fourier modes. In the isotropic strain component we find the strongest signal in zeroth and second mode, see SI Fig. 2 (B.i), which correspond to the constant and $\cos(2\varphi)$ angular Fourier modes. We then quantified the radial dependence of the second angular mode and we find that its magnitude decays as $r^{-2}$, as shown in Fig. 1 (C.ii). These results reminded us of the far field of Eshelby propagators in two-dimensional linear elastic sheets Eshelby (1957); Nicolas et al. (2018); Moriel et al. (2020) and we next tested whether linear elastic theory can indeed quantitatively account for our measurements.
We briefly present linear elastic theory of a sheet in which a point force dipole has been inserted. A force dipole is represented by a tensor $D_{ij}$ that can be decomposed into isotropic and traceless-symmetric contributions $D_{ij}=D_{0}\delta_{ij}+\tilde{D}_{ij}$  [^1] We consider only force dipoles that do not exert a net torque.. Inserting a force dipole in an elastic sheet induces a strain field that is determined by the force balance equation
$\partial_{j}\sigma_{ij}+\partial_{j}(\delta(\bm{r})D_{ij})$ $=0$ (1)
where $\sigma_{ij}$ is the stress tensor in the elastic sheet, which can be decomposed into pressure $P$ and shear stress $\tilde{\sigma}$: $\sigma_{ij}=-P\delta_{ij}+\tilde{\sigma}_{ij}$. The resulting strain field can be written as
$u_{ij}(\bm{r})=G_{ijkl}(\bm{r})D_{kl},$ (2)
where $G_{ijkl}$ is the elastic dipole propagator and $\bm{r}$ is the distance from the force dipole, see SI. We graphically illustrate the components of Eq. 2 in Fig. 1 (B.iii), using the Voigt tensor representation. In this representation, the force dipole propagator elements are organised in a matrix, which we plot in Fig. 1 (b). From this point, we use the normlised force dipole defined as $d_{ij}=D_{ij}/(2\pi K(1+\mu))$. The propagator elements depend only on $\cos(2\varphi)$, $\sin(2\varphi)$, $\cos(4\varphi)$ and $\sin(4\varphi)$, see Fig. 1 (B.iii) and SI, consistent with the observed second and fourth angular Fourier modes in the data. Furthermore, all components of $G_{ijkl}$ decay in space as $r^{-2}$. This is consistent with the the radial profile observed in the experimental data for $m=2$ modes, see Fig. 1 (C.ii) and SI Fig. 4. Radial profile of $m=4$ modes decay faster and appear to be much more susceptible to the noise in the data. Therefore, the strain field measured in the laser ablation experiments is consistent with a linear elastic response of the tissue to the force dipoles generated by the ablation. However, to confirm this hypothesis, we need to show that linear elastic theory can quantitatively account for the observed strain fields.
We fit the steady state tissue strain field $U^{ss}_{ij}$ with a strain field predicted by a linear elastic theory of a two-dimensional elastic sheet in Eq. 2. The fitting parameters are the three independent components of the force dipole tensor and the ratio $\mu$ of the two-dimensional shear and bulk elastic moduli. By analysing two laser ablation experiments, we find that the data are well described by the fit, as shown in Fig. 1 (D) and SI Fig. 5. The fitting parameters are consistent between the two experiments, see Table 1, except $\tilde{d}_{xx}$, which varies by a factor of two. This shows that there is a difference in the force dipole magnitude generated by the ablation, which could arise either from a difference of tissue pre-stresses or a difference in the effective ablation size in the two experiments. Furthermore, the ratio $\mu\approx 2$, corresponding to the Poisson ratio of $\nu\approx-0.33$, is comparable with the value $\mu\approx 3$, corresponding to the Poisson ration of $\nu\approx-0.5$, recently reported in the same tissue using a laser ablation method based on circular ablation shape Piscitello-Gómez et al. (2023).
To further test our analysis of the laser ablation experiments, we simulate a linear ablation in the vertex model of epithelial tissues Farhadifar et al. (2007). We remove two cells from a pre-stressed cellular network, see SI, and then we quantify the resulting strain. We find that the simulation shows the same pattern as in experiments, consistent with the linear elastic theory of a 2-d sheet Fig. 1 (E, F) and SI Fig. 6. We fit the linear elastic theory to the observed strain field pattern and infer the force density generated by the ablation from tissue deformation. Finally, we find the Poisson’s ratio of the vertex model cellular network to be $\nu=-0.061\pm 0.005$. This finding is consistent with the value predicted value $\nu\approx-0.060$ in Ref. Staddon et al. (2023) for the vertex model parameters we use, see SI.
Our results show that response of the D. melanogaster wing epithelium to a small linear laser ablation is consistent with that of a linearly elastic sheet with a negative two-dimensional Poisson ratio. This allows us to infer the force dipole tensor induced by the ablation from the tissue strain field, which we confirmed using the vertex model simulations.
Figure 1: (A.i) Fluorescently labeled cell outlines show the cellular packing before the laser ablation. The red bar indicates the location that will be ablated. (A.ii) Following the laser ablation tissue has deformed away from the ablated region. (A.iii) The cells we have tracked and segmented define the region of interest colored in white. (A.iv) In the region of interest, we show the accumulated cell strain component $U_{xy}$. The scale bar is $10\mu m$. (B.i) A force dipole in a two dimensional linear elastic system induces a strain in distance $\vec{r}$. (B.ii) Two dimensional strain field and force dipole tensors are decomposed into isotropic and traceless-symmetric components. (B.iii) A Voigt representation of the relation between strain at position $\vec{r}$ induced by a force dipole at the origin. (C.i)Angular Fourier modes of $\tilde{U}_{xy}$ at distance $r=16~{}\rm{\mu m}$, which corresponds to the yellow circle A.iv, shows a clear signal in modes $m=2$ and $m=4$ modes. (C.ii) The radial profile of the $m=2$ angular mode shows a radial profile consistent with $r^{-2}$ (dashed line). The dotted line indicates the distance $r=16~{}\rm{\mu m}$. (D.i) Fit of the linear elastic theory to the experimental data. The solid blue line is the fitted strain field with the corresponding fit parameters reported in Table I, exp. 1. (D.ii) Visualisation of the fitted strain field. (E.i) The cellular network of tissue in the vertex model simulation. The red bar indicates two cells removed by an ablation. (E.ii) The cell strain field accumulated in the tissue after the ablation. (F.i) Angular Fourier modes of the strain field pattern in E.ii showing peaks in the second and fourth modes. (F.ii) Fit of the linear elastic theory to cell strain field in the vertex model. The solid blue is the fitted strain field with the fit parameters $d_{0}=(10.92\pm 0.01)\times 10^{-2}a_{0}$, $\tilde{d}_{xx}=(8.29\pm 0.01)\times 10^{-2}a_{0}$, $\tilde{d}_{xy}=(2.93\pm 0.01)\times 10^{-2}a_{0}$ and $\mu=1.13\pm 0.01$.
II.2 Viscous dissipation determines tissue relaxation dynamics
Our analysis of the strain in the ablated tissue after it has relaxed allowed us to study its elastic properties. Now, we explore dynamics of tissue strain as it approaches the equilibrium, which will allow us to study dissipative processes in fly wing epithelium. We consider two possible dissipation mechanisms: viscosity and friction with a substrate.
Interestingly, in a continuum model of a tissue as an elastic sheet, these two dissipation mechanisms lead to qualitatively different relaxation dynamics upon insertion of a force dipole. In particular, in an elastic sheet with only viscous dissipation, all strain field components decays as $1/r^{2}$ with the distance from the force dipole throughout the relaxation with the amplitude evolving in time, see SI and Fig. 2 (A.i). However, in an elastic sheet with only frictional dissipation, the strain field propagates through the sheet: at distances smaller than the propagation front, the strain field converges to the steady state solution and rapidly decays beyond the front, see SI and Fig. 2 (A.ii).
To quantify the dynamics of tissue relaxation following the laser ablation, we measure the strain field in time intervals of $4.5s$, starting from $1s$ after the ablation, (SI Fig. 7). We find that the angular strain pattern is the same as in the mechanical equilibrium one, see Movie 1. We analyse the dynamics of the strain radial profile by calculating the second angular Fourier mode of the strain field at all measured timepoints, see Fig. 2 (B.i) and SI Fig. 7. We find that at all timepoints the strain field is consistent with $1/r^{2}$ decay and we do not see any signature of a propagating front, see Fig. 2. Therefore, we conclude that the viscosity is the dominant mode of dissipation in the fly wing epithelium on timescales above $1s$, corresponding to the initial timepoint we observe after the ablation, and below $T_{e}\approx 50s$.
To make sure that a predicted propagation front could be observed in presence of frictional dissipation we performed vertex model simulations of laser ablation that include only the frictional dissipation, see SI for details. We indeed find a clear propagation of the strain field that converges to the steady state value as the propagation front expands (2 (B.ii) and SI Fig. 8), as predicted by the analystical calculation (2 (A.ii)).
Figure 2: Time evolution the second Fourier mode $m=2$ of the strain field $xy$ component induced by a force dipole for: (A.i) linear elastic sheet with viscous dissipation, (A.ii) linear elastic sheet with frictional dissipation, (B.i) laser ablation experiment (B.ii) vertex model simulation with frictional dissipation. Dashed line indicates the power-law $r^{-2}$ decay for reference.
Table 1: Parameters obtained by fitting the continuum theory to spatio-temporal pattern of strain field accumulated in the tissue after the ablation.
exp. $d_{0}~{}[\mu m]^{2}$ $\tilde{d}_{xx}~{}[\mu m]^{2}$ $\tilde{d}_{xy}~{}[\mu m]^{2}$ $\mu$ $\tau_{1}~{}(s)$ $\tau_{2}~{}(s)$
1 $16\pm 1$ $15\pm 1$ $2\pm 1$ $1.9\pm 0.1$ $11\pm 1$ $12\pm 1$
2 $15\pm 2$ $31\pm 2$ $0\pm 2$ $2.1\pm 0.3$ $48\pm 2$ $55\pm 2$
Our continuum model of an elastic sheet provides the dynamical strain propagator ${G}_{ij}(\bm{r},t)={G}_{ij}(\bm{r}){T}_{ij}(t)$, where no sum over repeated indices is performed. ${G}_{ij}(\bm{r})$ is the steady state propagator, and ${T}_{ij}(t)$ is the relaxation matrix whose components contains two relaxation factors $g_{1}(t)=1-\exp{(-t/\tau_{1})}$ and $g_{2}(t)=1-\exp{(-t/\tau_{2})}$, where the two relaxation timescales
$\tau_{1}$ $=\frac{\overline{\eta}}{\overline{K}}\frac{1+2\eta/\overline{\eta}}{1+\mu}$ (3)
$\tau_{2}$ $=\frac{\eta}{K}$ (4)
are related to the bulk and shear elastic constants, and bulk and shear viscosities $\overline{K},2K,\overline{\eta},\text{ and }2\eta$, respectively (see SI). To extract $\tau_{1}$ and $\tau_{2}$ from experimental data, as described in SI and shown in SI Fig. 1. We find that the two relaxation timescales are very similar, although they vary between experiments, see Table 1. It is at the moment unclear whether the difference between the relaxation timescales in two experiments stems from variability of elastic constants or viscosity, or some other factor. While this will be an important open question to resolve in future, further investigation of relaxation timescale variability is beyond the scope of this work.
So far, our detailed analysis of the strain in the fly wing epithelium in response to the laser ablation shows that the tissue behaves as a two-dimensional viscoelastic solid. With this knowledge in hand, we can now infer force dipoles generated by cell divisions, and thereby study their mechanics, by simply observing the dynamics of the surrounding tissue strain caused by the division, without the need to directly perturb the tissue.
III Elastic deformation imprinted by a cell division
Cell division is a complex process during which the mother cell is split into two daughter cells. Here, we aim to characterise the mechanical forces generated by the dividing cell by analysing the strain field a dividing cell induces in the fly wing epithelium.
We analysed cell divisions in previously published experiments Dye et al. (2021) on explanted D. melanogaster wing imaginal discs  [^2] Note that this developmental stage preceeds the pupal stage in which linear laser ablations were performed Piscitello-Gómez et al. (2023) and although both are fruit fly wing tissues, it is possible that tissue elastic constants have changed between the two stages.. Wing discs were cultured over about $13$ hours, during which the tissue was imaged every $5$ minutes, and individual cells were segmented and tracked over time. The progress of a typical cell division is shown in Fig. 3 (A). Before the actual cell division, the mother appears to inflate Fig. 3 (A.ii), which is not a real increase in cell size, but is rather a consequence of the cell nucleus arriving to the apical surface of the cell to divide Gibson et al. (2011); Strzyz et al. (2016). This transient inflation is followed by the creation of the new cell boundary, and the two daughter cells become visible Fig. 3 (A.iii). For each cell division, we define the time of the new bond creation as the reference cell division time $t=0$.
Fig. 3 (A.iv) qualitatively summarizes these experimental observations. For each cell division, we quantify the tissue strain field in the vicinity of the dividing cell, as described in Section II, relative to the starting timepoint $t_{0}=-60~{}\rm{min}$ before the cell division time. We show an example of the cell strain field induced by a dividing cell in Fig. 3 (B). The timescale on which the cell division forces evolve, which is on the order of tens of minutes, is much larger than the relaxation timescale extracted in Section II, which is below one minute. Therefore, the observed strain field dynamics is largely due to changes of the force dipole generated by the cell division, and not due to much faster stress propagation through the tissue.
Figure 3: (A) Snapshots of a cell division in three stages. (A.i) As the mother cell starts the division process, (A.ii) it rounds up, corresponding to the inflation, then (A.iii) it shrinks and divides into the daughter cells. The red dots show the mother and the daughter cells. (A.iv) A schematic representation of the cell area of the mother cell (brown) and the daughter cells (orange) during the division process. The symbols highlight the stage of the dividing cell before the division, at the peak of inflation, and after the division. The new bond forms at $t=0$, and $t_{0}$ is the reference state. (B) Strain $\tilde{U}_{xy}$ of the cells near a typical cell division at the three stages shown in (A) with $t_{0}=-60~{}[\rm{min}]$. (C) Ensemble average ($N=682$) of $\tilde{U}_{xy}$ around the dividing cell: (C.i) at the peak of the inflation $t=-20~{}[\rm{min}]$ (corresponding to (A.ii) and (B.ii)), and (C.ii) after the division $t=30~{}[\rm{min}]$ (corresponding to (A.iii) and (B.iii)). (D) The force dipole generated by a cell division as a function of time. (E) A schematic summary of the force dipoles during stages of a cell division.
We find that a single cell division generates weaker forces compared to those induced by laser ablation (compare Fig. 1 A and Fig. 3 B), and the corresponding strain field is more susceptible to noise. Therefore, we average the strain field from multiple cell divisions ($N=682$). We first align the centers of the cell division, defined as the area weighted average of daughter cell centers at the cell division time ($t=0$). We next rotate each division such that the orientation of the line connecting daughter cell centers is oriented along the $x-$axis at the cell division time. We show the resulting strain field component $\tilde{U}_{xy}$ in Fig. 3 (C) at times $t=-20~{}\rm{min}$, corresponding to the peak of the cell inflation, and $t=30~{}\rm{min}$ after the cell division. In Fig. 3 (C.i), the 2-fold symmetry of the angular strain pattern indicates that the isotropic force dipole is dominant, consistent with the isotropic cell inflation. In Fig. 3 (C.ii), the pattern is closer to 4-fold symmetric one, such as would be induced by an anisotropic force dipole.
We fit the linear elastic theory Eq. 2 to the experimentally measured strain field at each time point and extract the resulting force dipole $d_{ij}(t)$. We find that the isotropic force dipole component $d_{0}$ rises sharply during the mother cell inflation and peaks at about $t=-20~{}\rm{min}$ (magenta line in Fig. 3 D). However, by the time of the new bond formation, the isotropic component has disappeared and remains negligible at later times. This result shows that cell does not noticeably grow during the division, but rather that the division reorganises the available material into the two daughter cells. The traceless part of the force dipole $\tilde{d}_{xx}$ grows as the new bond is formed and persists in the tissue multiple hours after the division (green line in Fig. 3 D), indicating that the division-induced strains remain imprinted in the tissue long after the division has been executed. Over time, $\tilde{d}_{xx}$ slowly decays while its fluctuations increase, representing the effect of other cell divisions and rearrangements that appear in the tissue. We estimate the tissue fluidization time from the decay of $\tilde{d}_{xx}$ in time by fitting an exponential function to it. We find the fluidization timescale to be $\tau_{\alpha}\simeq 3.5\pm 0.7$ minutes, over which $\tilde{d}_{xx}$ vanishes and, therefore, the strain field induced by the division is erased. Note that component $\tilde{d}_{xy}$ remains negligible at all times.
Finally, we note that at the time of the peak of $d_{0}$, there is a small but clear negative peak in $\tilde{d}_{xx}$. To interpret this observation, we have to remember that our strain measurements are relative to the state of the cells at $t_{0}=-60~{}\rm{min}$. Since the mother cell attains an almost perfectly circular shape during the inflation, any initial anisotropy of the mother cell at time $t_{0}$ gives rise to an anisotropic force dipole component during the inflation. Therefore, the observed negative value of $\tilde{d}_{xx}$ indicates that mother cells are slightly elongated along the future division axis, consistent with the well known Hertwig’s rule Hertwig (1884); Théry et al. (2007); Gibson et al. (2011); Lisica et al. (2022); Middelkoop et al. (2023). We summarise these results in Fig. 3 (E).
IV Discussion
In this work, we devised a full spatio-temporal analysis of tissue strain induced by two types of force dipole fields: linear laser ablations and cell divisions. In this way, we have shown that the developing Drosophila wing epithelium behaves as a two-dimensional elastic solid with viscous dissipation. We also measured its two-dimensional Poisson ratio, which we find to be negative, consistent with the values reported in the literature Dye et al. (2021); Piscitello-Gómez et al. (2023).
We characterized the dynamics of mechanical forces generated by a dividing cell throughout the cell division process. We found that strain fields imprinted by a cell division are slowly erased on the timescale of $\tau_{\alpha}=3.5\pm 0.7~{}\rm{hr}$, which can be understood as a fluidization timescale of the Drosophila wing tissue. Furthermore, our analysis shows that there is no cell growth during a division and stresses induced by the cell division correspond to those of a pure shear force dipole. Finally, we found that Hertwig’s rule applies in the D. melanogaster wing disc Gibson et al. (2011); Mao et al. (2013).
An important implication of our work is that the Drosophila wing epithelium is a solid whose fluidization time $\tau_{\alpha}$ is not negligible on the developmental timescales. For example, large scale shape changes that the wing undergoes during eversion are executed on a similar timescale Fuhrmann et al. (2023). A possible physical picture of the Drosophila wing epithelium that could account for our observations, is that of a glassy material, or as was put forward in glass literature, a solid that flows Dyre (2023); Ozawa and Biroli (2023); Tahaei et al. (2023). In other words, the long-range elastic interactions between cells would lead to correlated glassy dynamics, which has been reported in tissue cultures Angelini et al. (2011); Park et al. (2015), and it remains to be seen whether they indeed play a role in developing tissues. In summary, our results quantify force dipoles generated by dividing cells in developing D. melanogaster wing epithelia, as well es elastic interactions between cells, which paves the way to development tissue mechanics theory and motivates further experimental study of collective cell dynamics.
Acknowledgements
We would like to thank Michael Staddon, Frank Jülicher, Jonas Neipel for usefull discussions and comments. NAD, JFF, and GC acknowledge funding from the Max Planck Society, Excellence Cluster Physics of Life (EXC-2068-390729961), Deutsche Forschungsgemeinschaft (SPP1782/EA4/10-2), Deutsche Krebshilfe/MSNZ Dresden.
V Methods
V.1 Cell segmentation and tracking
The linear ablation experiments came from the pupal morphogenesis stage and were previously published in Ref. Piscitello-Gómez et al. (2023), where the authors analyzed initial retraction velocity of the ablated bonds. Here, we took these experiments and fully segmented and tracked cells in the entire field using the FIJI plugin, TissueAnalyzer Aigouy et al. (2016), using a bond cutoff of 2 pixels. Then, after the manual correction of segmentation errors, we use TissueMiner Etournay et al. (2016) to analyse changes in the cell shape and size.
The cell division experimental data came from the wing disc stage and was previously segmented, tracked and analyzed in Ref. Dye et al. (2017, 2021).
V.2 Tissue relaxation after a laser ablation
After the ablation, nearby tissue relaxes to the new mechanical equilibrium. However, beyond $2$ minutes after the ablation we observed additional motion of the tissue which we interpret as the onset of the would healing response in the tissue. In this study, we are interested in the stress released by the ablation in the tissue, not the wound healing process, and therefore we limit our analysis to times up to $2$ minutes after the laser ablation.
V.3 Tissue strain field measurement
Strain field consists of isotropic strain $U_{kk}(\mathbf{r},t)/2$ and shear strain $\tilde{U}_{ij}(\mathbf{r},t)$. To measure the strain field in the tissue we calculate the change in cell shape and size as described in Merkel et al. (2017); Duclut et al. (2021) at each timepoint. Since there are no cell rearrangements and divisions we estimate the accumulated tissue shear strain field in a cell from the difference between the cell shape tensors
$\tilde{U}_{ij}(t)$ $\simeq Q_{ij}(t)-Q_{ij}(0)$ (5)
The only approximation is that we neglect to account for change in cell elongation due to local rotations. We determined the exact cell shear tensor at several timepoints and we found that the difference from our approximation are small. Therefore, we use the approximate expression for simplicity. We estimate the isotropic strain field in a cell from the change in cell area $a$
$U_{kk}(t)$ $=\log{\left(\frac{a(t)}{a(0)}\right)}\quad.$ (6)
We define the tissue strain field $U_{ij}(\mathbf{r},t)$ at position $\mathbf{r}$ to have the value of the strain tensor obtained for the cell inside of which point $\mathbf{r}$ lies.
V.4 Discrete angular Fourier transform
To investigate the angular pattern of the strain field and study its radial profiles, we look at the angular discrete Fourier transformation of the tissue strain field. The strain in position $\bm{r}$ is given by ${U}_{ij}(r,\varphi)$ where $r$ is the distance to the center of the cut and $\varphi$ is the polar angle. We use the notation ${U}_{ij}(r,m)$ for the $m$-th angular mode of discrete angular Fourier transform at a distance $r$, and it is defined as,
${U}_{ij}(r,m)=\frac{1}{N}\sum_{i=1}^{N}{U}_{ij}(r,\varphi_{i})e^{im\varphi_{i}},$ (7)
where $\varphi_{i}$s are $N$ uniformly distributed number between 00 and $2\pi$.
V.5 Fit of the elastic theory to data
We measure force dipole, the ratio of shear modules to bulk modules, and viscosity by fitting the strain field pattern to the elastic and viscoelastic theories. First, we define a cost function
$E=\sum_{\alpha\in\rm{cells}}\Big{[}(U_{kk}^{\alpha}-u_{kk}(\bm{r}^{\alpha})^{2}+(\tilde{U}_{xx}^{\alpha}-\tilde{u}_{xx}(\bm{r}^{\alpha})^{2}+\\ (\tilde{U}_{xy}^{\alpha}-\tilde{u}_{xy}(\bm{r}^{\alpha})^{2}\Big{]},$ (8)
where ${U}_{ij}^{\alpha}$ is the cell strain measured from data, $\bm{r}^{\alpha}$ is the position of centre of each cell and ${u}_{ij}(\bm{r}^{\alpha})$ is the strain field predicted from theory in position $\bm{r}^{\alpha}$. By minimising $E$, we find the force magnitude of the force dipole, the ratio of the shear modulus to bulk modulus, and the relaxation timescales. In the elastic case, first, we minimise a steady state signal to find $\mu$ then we fix $\mu$ and repeat the minimisation process to find ${d}_{ij}$. In viscoelastic case, we use $\mu$ and ${d}_{ij}$ from the steady state elastic response, then we minimise $E$ at different time points to find $T_{ij}(t)$.
To analyse cell divisions, we first choose timepoint $t=25~{}\rm{min}$ and change the reference time $t_{0}$ from $-120$ to $-40$. We fit the data for different $t_{0}$ in that range and obtain corresponding set of $\mu(t_{0})$ values. We determine the parameter $\mu$ as the average $\langle\mu(t_{0})\rangle$ and we do not vary it in subsequent fitting of the strain field. Then, to determine the force dipole at each timepoint $t$ we fix $t_{0}=-60~{}\rm{min}$ fit the theory to the data to find the force dipole at different time points.
V.6 Estimation of the tissue fluidization time
To measure the decay time scale of the anisotropic force dipole $\tilde{d}_{xx}$, we fit an exponential function to normalized force dipole $y=\tilde{d}_{xx}(t)/\langle a\rangle$ as $f(y)=\tilde{d}_{xx}(t=0)\exp[-t/\tau_{\alpha}]$ to observed data in Fig. 3 (D). From that, we obtain the fluidization times scale $\tau_{\alpha}=3.5\pm 0.7\rm{hr}$ and $\tilde{d}_{xx}(t=0)/\langle a\rangle=0.041\pm 0.001$.

### v2

@@
 Therefore, a quantitative description of these force dipoles, and their consequences on tissue mechanics, is one of the central problems in understanding the overall tissue mechanics.
-In this work we analyze previously published experimental data on fruit fly D.
+In this work, we analyze previously published experimental data on fruit fly D.
 melanogaster wing epithelia to quantitatively describe the deformation fields induced by a cell-scale force dipole.
-We find that the measured deformation field can be explained by a simple model of fly epithelium as a linearly elastic sheet.
-This fact allows us to use measurements of the strain field around cellular events, such as cell divisions, to infer the magnitude and dynamics of the mechanical forces they generate.
-In particular, we find that cell divisions exert a transient isotropic force dipole field, corresponding to the temporary localisation of the cell nucleus to the tissue surface during the division, and traceless-symmetric force dipole field that remains detectable from the tissue strain field for up to about $3.5$ hours after the division.
+We find that the measured deformation field can be explained by a simple model of wing epithelium as a linearly elastic sheet.
+This allows us to infer the magnitude and dynamics of the mechanical forces generated by the cell divisions.
+In particular, we find that cell divisions exert a transient isotropic force dipole field, corresponding to the temporary localization of the cell nucleus to the tissue surface during the division, and traceless-symmetric force dipole field that remains detectable from the tissue strain field for up to about $3.5$ hours after the division.
 This is the timescale on which elastic strains are erased by other mechanical processes and therefore it corresponds to the tissue fluidization timescale.
 In summary, we have developed a method to infer force dipoles induced by cell divisions, by observing the strain field in the surrounding tissues.
-Using this method we quantitatively characterize mechanical forces generated during a cell division, and their effects on the tissue mechanics.
+Using this method we quantitatively characterize mechanical forces generated during a cell division and their effects on the tissue mechanics.
 I Introduction
@@
 (2010) that are elastic on sufficiently short timescales.
-Tissue fluidization on longer timescales arises from cell divisions, extrusions and intercalations, that restructure the tissue over time Ranft et al.
+Tissue fluidization on longer timescales arises from cell divisions, extrusions, and intercalations, which restructure the tissue over time Ranft et al.
 (2010); Etournay et al.
@@
 (2021).
-Cell divisions are an intrinsic source of mechanical force generation within tissues.
-When a cell divides, particularly in a closely packed epithelium, it induces deformation in the surrounding tissue Gibson et al.
+Cellular processes, such as cell divisions and cell extrusions, are intrinsic sources of mechanical force generation within tissues, inducing deformation in the surrounding tissue Gibson et al.
 (2011); Mao et al.
-(2013), potentially leading to cellular rearrangements Firmino et al.
-(2016) and contributing to the overall tissue flow Ranft et al.
+(2013); Roellig et al.
+(2022), which can in turn lead to cellular rearrangements Firmino et al.
+(2016) and contribute to the overall tissue flow Ranft et al.
 (2010); Matoz-Fernandez et al.
@@
 (2013).
-However, despite the fact that cell divisions are a hallmark of biological matter and seem to be major factor in controlling material properties of biological tissues, quantitative understanding of the mechanical forces generated by cell divisions and their immediate consequences on the surrounding tissue is still lacking.
-In order to study the mechanics of cell divisions, we need to probe mechanical stresses in the surrounding tissue.
+However, despite the fact that cell divisions are a hallmark of biological matter and seem to be a major factor in controlling material properties of biological tissues, a quantitative understanding of the mechanical forces generated by cell divisions and their immediate consequences on the surrounding tissue is still lacking.
+To study the mechanics of cell divisions, we need to probe mechanical stresses in the surrounding tissue.
 One of the most common methods of probing mechanical stresses is the laser ablation experiment, in which one or multiple cells are destroyed by a laser.
@@
 (2023).
-Information about the pre-stresses are extracted either from the initial dynamics of tissue boundary displacement Mayer et al.
+Pre-stresses are then characterised either by the initial dynamics of tissue boundary displacement Mayer et al.
 (2010); Kasza et al.
@@
 (2023).
-Previous analyses focused on changes in position or shape of specific points or outlines in the vicinity of the ablation.
+Previous analyses focused on changes in the position or shape of specific points or outlines in the vicinity of the ablation.
 Therefore, they could not provide an insight into the strain field induced in the surrounding tissue, and they had to rely on assumptions about tissue mechanical properties.
@@
 If we could instead infer force dipoles induced by cellular processes such as cell divisions, we would be able to extract information about tissue mechanics in a completely non-invasive way.
-Such a method was recently developed to infer the properties of the core of plastic particle rearrangements in computer glasses Moriel et al.
-(2020, 2024).
-Therefore, the starting point of our work, on the way to study cell division mechanics, is to develop a method that allows us infer the force dipole in an epithelial tissue from the tissue strain field.
+Therefore, the starting point of our work, on the way to study cell division mechanics, is to develop a method that allows us to infer the force dipole in an epithelial tissue from the tissue strain field.
+The idea to infer properties of a localized force field from the strain field it induces was recently employed to characterize the properties of the core of plastic events in computer glasses Moriel et al.
+(2020); Chacko et al.
+(2021); Moriel et al.
+(2024).
+Here, we proceed independently of these results using an approach better suited for epithelial tissue data.
 In this work, we first provide a detailed quantitative analysis of the tissue displacement field following a linear laser ablation performed in the pupal wing epithelium of the fruit fly D.
@@
 This allows us to infer the force dipole induced by the ablation, normalized by the elastic constant, from the observed strain field.
-Furthermore, by analyzing the dynamics of the strain field, we show that the dominant mode of dissipation in the fly wing epithelium on short timescales of $1s$ to $50s$ is the viscous dissipation, as opposed to frictional dissipation with a substrate.
+Furthermore, by analyzing the dynamics of the strain field, we show that the dominant mode of dissipation in the fly wing epithelium on short timescales less than a minute is viscous dissipation, as opposed to frictional dissipation with a substrate.
 Using our method to infer force dipole from its strain field, we determine the force dipole generated by cell divisions in the D.
@@
 Furthermore, we find that the traceless-symmetric component of the force dipole, imprinted by the cell division, remains visible in the surrounding tissue for up to about $3.5$ hours after the division.
-This provides a measurement of the fluidization timescale in the tissue, and demonstrates importance of elastic interactions in tissue-scale morphogenetic processes that take place on a comparable timescale, such as wing disc eversion Fuhrmann et al.
+This provides a measurement of the fluidization timescale in the tissue and demonstrates the importance of elastic interactions in tissue-scale morphogenetic processes that take place on a comparable timescale, such as wing disc eversion Fuhrmann et al.
 (2023).
@@
 II.1 Linear laser ablation experiments
-Linear laser ablation experiments, during which a small linear segment of tissue is destroyed by a laser, have been performed in the D.
-melanogaster wing epithelium during pupal morphogenesis Piscitello-Gómez et al.
+In linear laser ablation experiments, a small linear segment of tissue is destroyed by a laser.
+In Ref.
+Piscitello-Gómez et al.
+(2023), laser ablation experiments have been performed in the D.
+melanogaster wing epithelium during pupal morphogenesis.
+Information about the pre-stresses in the tissue was inferred from the dynamics of the nearest cell-cell junction as it retracted away from the ablation.
+Here, we use laser ablation experiment data published in Ref.
+Piscitello-Gómez et al.
+(2023) to study the deformation field in the tissue surrounding the laser ablation.
+An image of a wing tissue before and after such a laser ablation experiment is shown in Fig.
+1 (A.i), where the red line indicates the line ablated by the laser.
+Ablations were performed in a plane thinner than $1~{}\mu m$ on the apical surface of the tissue Piscitello-Gómez et al.
 (2023).
-To obtain information about the pre-stress in the tissue, the dynamics of the closest cell-cell junction was measured as it retracted away from the ablation.
-Here, we use movies of laser ablation experiments published in Ref.
+Following the ablation, the tissue relaxes until reaching a mechanical equilibrium after about $T_{e}\approx 30s$ after ablation.
+We show an example of the relaxed state in Fig.
+1 (A.ii).
+The tissue remains in this state until after around $1\rm{min}$.
+After this time, the tissue starts to exhibit additional flows, directed towards the line of ablation, which likely correspond to the wound-healing response of the tissue, see Methods and SI.
+Note that in most experiments from Ref.
 Piscitello-Gómez et al.
-(2023) to study the deformation in the tissue surrounding the laser ablation.
-An example of such a laser ablation experiment is shown in Fig.
-1 (A.i), where the red line indicates the line ablated by the laser.
-Following the ablation, the tissue relaxes until reaching a mechanical equilibrium after at most $T_{e}=50s$ in Fig.
-1 (A.ii), see Methods.
-To measure the tissue strain field, we segmented the apical cell surfaces over time and quantified the area $a$ and the elongation tensor $Q_{ij}$ of each cell, which quantifies its shape Etournay et al.
+(2023) we could not segment and analyze a reasonably large circle of cells around the ablation (SI Fig.
+1), we selected four experiments that we could best analyze.
+In the main text we present analysis of the clearest one, denoted Experiment 1.
+In Experiment 2, such flows appear earlier so that no intermediate relaxed state can be identified.
+In Experiments 3 and 4 the strain field generated by the ablation is also weaker than in Experiment 1, so that the noise overcomes the signal already at small distances from the ablation.
+Analyses of these three experiments are presented in the SI.
+To measure the tissue strain field, we segment the apical cell surfaces over time and quantify area $a$ and the elongation tensor $Q_{ij}$ of each cell.
+The cell elongation tensor is defined such that the co-rotational changes in the elongation tensor $Q_{ij}$ correspond to the accumulated pure shear strain Etournay et al.
 (2015); Merkel et al.
@@
 In Fig.
-1 (A.iv) we show the pure shear strain field component $\tilde{U}_{xy}$ at $100s$ after the laser ablation, which exhibits a clear spatial pattern, see SI Fig.
-2 (A) for the other traceless-symmetric component and isotropic component of experiment 1 and SI Fig.
-3 (A) for all components of the experiment 2.
-Our goal is to find the relationships between the induced strain field and the forces induced by the ablation.
+1 (A.iv) we show the pure shear strain field component $\tilde{U}_{xy}$ at $50s$ after the laser ablation, which exhibits a clear spatial pattern.
+See SI Fig.
+3 (A) for the other traceless-symmetric component and the isotropic component.
+Our goal is to find the relationships between the induced strain field and the stresses induced by the ablation.
 Since the process of ablation does not introduce any net force, the dominant component of the induced force field will be dipolar.
@@
 1 (B.i,ii).
-We first consider tissue strain after it has reached the mechanical equilibrium after the time $T_{e}$ and calculate the discrete angular Fourier spectrum $\tilde{U}^{ss}_{ij}(r,m)$, where and $r=|\bm{r}|$ is distance from the center of ablation and $m$ is the angular mode index, see Methods.
+We first consider tissue strain after it has reached the mechanical equilibrium after the time $T_{e}$ and calculate the discrete angular Fourier spectrum $\tilde{U}^{ss}_{ij}(r,m)$, where $r=|\bm{r}|$ is the distance from the center of ablation and $m$ is the angular mode index, see Methods.
 We find the strongest signal in the second and the fourth mode of the pure shear strain components, see Fig.
@@
 3 (B.ii), which correspond to $\cos(2\varphi)$ and $\cos(4\varphi)$ angular Fourier modes.
-In the isotropic strain component we find the strongest signal in zeroth and second mode, see SI Fig.
+In the isotropic strain component, we find the strongest signal in the zeroth and second mode, see SI Fig.
 3 (B.i), which correspond to the constant and $\cos(2\varphi)$ angular Fourier modes.
-We then quantified the radial dependence of the second angular mode and we find that its magnitude decays as $r^{-2}$, as shown in Fig.
+We then quantify the radial dependence of the second angular mode and we find that its radial profile is consistent with $r^{-2}$, as shown in Fig.
 1 (C.ii).
@@
 (2018); Moriel et al.
-(2020) and we next tested whether linear elastic theory can indeed quantitatively account for our measurements.
-We briefly present linear elastic theory of a sheet in which a point force dipole has been inserted.
+(2020), and we next tested whether the linear elastic theory can indeed quantitatively account for our measurements.
+We briefly present the linear elastic theory of a sheet in which a point force dipole has been inserted.
 A force dipole is represented by a tensor $D_{ij}$ that can be decomposed into isotropic and traceless-symmetric contributions $D_{ij}=D_{0}\delta_{ij}+\tilde{D}_{ij}$  [^1] We consider only force dipoles that do not exert a net torque..
@@
 1 (B.iii), using the Voigt tensor representation.
-In this representation, the force dipole propagator elements are organised in a matrix, which we plot in Fig.
+In this representation, the force dipole propagator elements are organized in a matrix, which we plot in Fig.
 1 (B).
-From this point, we use the normlised force dipole defined as $d_{ij}=D_{ij}/(2\pi K(1+\mu))$.
+From this point, we use the normalized force dipole defined as $d_{ij}=D_{ij}/2K$.
 The propagator elements depend only on $\cos(2\varphi)$, $\sin(2\varphi)$, $\cos(4\varphi)$ and $\sin(4\varphi)$, see Fig.
@@
 1 (C.ii) and SI Fig.
-4.
-Radial profile of $m=4$ modes decay faster and appear to be much more susceptible to the noise in the data.
+5.
+The radial profiles of $m=4$ modes are more noisy so that it is difficult to characterise the radial decay.
 Therefore, the strain field measured in the laser ablation experiments is consistent with a linear elastic response of the tissue to the force dipoles generated by the ablation.
@@
 The fitting parameters are the three independent components of the force dipole tensor and the ratio $\mu$ of the two-dimensional shear and bulk elastic moduli.
-By analysing two laser ablation experiments, we find that the data are well described by the fit, as shown in Fig.
+By analyzing the ablation experiment, we find that the data are well described by the fit, as shown in Fig.
 1 (D) and SI Fig.
-5.
-The fitting parameters are consistent between the two experiments, see Table 1, except $\tilde{d}_{xx}$, which varies by a factor of two.
-This shows that there is a difference in the force dipole magnitude generated by the ablation, which could arise either from a difference of tissue pre-stresses or a difference in the effective ablation size in the two experiments.
-Furthermore, the ratio $\mu\approx 2$, corresponding to the Poisson ratio of $\nu\approx-0.33$, is comparable with the value $\mu\approx 3$, corresponding to the Poisson ration of $\nu\approx-0.5$, recently reported in the same tissue using a laser ablation method based on circular ablation shape Piscitello-Gómez et al.
+6 (A).
+The fitted parameter values are: $d_{0}=160\pm 10~{}\mu m^{2}$, $\tilde{d}_{xx}=150\pm 10~{}\mu m^{2}$, $\tilde{d}_{xy}=20\pm 10~{}\mu m^{2}$, $\mu=1.9\pm 0.1$, see Methods.
+Furthermore, the ratio $\mu\approx 2$, corresponding to the Poisson ratio of $\nu\approx-0.33$, is comparable with the value $\mu\approx 3$, corresponding to the Poisson ratio of $\nu\approx-0.5$, recently reported in the same tissue using a laser ablation method based on circular ablation shape Piscitello-Gómez et al.
 (2023).
@@
 (2007).
-We remove two cells from a pre-stressed cellular network, see SI, and then we quantify the resulting strain.
-We find that the simulation shows the same pattern as in experiments, consistent with the linear elastic theory of a 2-d sheet Fig.
+We remove two cells from a pre-stressed cellular network, where the pre-stress represents the proximal-distal stress in the wing epithelium Etournay et al.
+(2015), and the ablation axis is perpendicular to the pre-stress orientation, recapitulating the orientation of ablation in experiments, see SI for details.
+We then quantify the resulting strain and find that the simulation shows the same pattern as in experiments, consistent with the linear elastic theory of a 2D sheet Fig.
 1 (E, F) and SI Fig.
-6.
+7.
 We fit the linear elastic theory to the observed strain field pattern and infer the force density generated by the ablation from tissue deformation.
-Finally, we find the Poisson’s ratio of the vertex model cellular network to be $\nu=-0.061\pm 0.005$.
-This finding is consistent with the value predicted value $\nu\approx-0.060$ in Ref.
-Staddon et al.
-(2023) for the vertex model parameters we use, see SI.
-Our results show that response of the D.
+Finally, we find the Poisson’s ratio of the vertex model cellular network to be $\nu=-0.13\pm 0.04$.
+This finding is consistent with the value $\nu=-0.095\pm 0.009$ found by a direct numerical measurement of the shear and bulk modulus of the same network, see SI.
+Our results show that the response of the D.
 melanogaster wing epithelium to a small linear laser ablation is consistent with that of a linearly elastic sheet with a negative two-dimensional Poisson ratio.
@@
 The red bar indicates the location that will be ablated.
-(A.ii) Following the laser ablation tissue has deformed away from the ablated region.
-(A.iii) The cells we have tracked and segmented define the region of interest colored in white.
-(A.iv) In the region of interest, we show the accumulated cell strain component $U_{xy}$.
+(A.ii) Cell outlines $50s$ after the laser ablation.
+(A.iii) Tracked and segmented cells in the region of interest are colored in white.
+(A.iv) The accumulated cell strain component $U_{xy}$, $50s$ after the laser ablation.
 The scale bar is $10\mu m$.
-(B.i) A force dipole in a two dimensional linear elastic system induces a strain in distance $\vec{r}$.
-(B.ii) Two dimensional strain field and force dipole tensors are decomposed into isotropic and traceless-symmetric components.
+(B.i) A force dipole in a two-dimensional linear elastic system induces a strain in distance $\vec{r}$.
+(B.ii) Decomposition of two-dimensional strain field and force dipole tensors into isotropic and traceless-symmetric components.
 (B.iii) A Voigt representation of the relation between strain at position $\vec{r}$ induced by a force dipole at the origin.
-(C.i)Angular Fourier modes of $\tilde{U}_{xy}$ at distance $r=16~{}\rm{\mu m}$, which corresponds to the yellow circle A.iv, shows a clear signal in modes $m=2$ and $m=4$ modes.
+(C.i) Angular Fourier modes of $\tilde{U}_{xy}$ at distance $r=16~{}\rm{\mu m}$, which corresponds to the yellow circle A.iv, show a clear signal in modes $m=2$ and $m=4$.
 (C.ii) The radial profile of the $m=2$ angular mode shows a radial profile consistent with $r^{-2}$ (dashed line).
 The dotted line indicates the distance $r=16~{}\rm{\mu m}$.
-(D.i) Fit of the linear elastic theory to the experimental data.
-The solid blue line is the fitted strain field with the corresponding fit parameters reported in Table I, exp.
-1.
-(D.ii) Visualisation of the fitted strain field.
-(E.i) The cellular network of tissue in the vertex model simulation.
-The red bar indicates two cells removed by an ablation.
-(E.ii) The cell strain field accumulated in the tissue after the ablation.
+Analysis in both C.i and C.ii is performed on the strain field averaged over the time interval from $30s$ to $70s$ after laser ablation and the error bars show the standard deviation over this interval.
+(D.i) The fit of the linear elastic theory to the experimental data, shown in C.
+The solid blue line is the fitted strain field with the corresponding fit parameters reported in the text.
+(D.ii) Visualization of the fitted strain field.
+(E.i) The cellular network in the vertex model simulation.
+The red bar indicates two cells removed by ablation.
+(E.ii) The cell strain field accumulated in the tissue after reaching the steady state following the simulated ablation.
 (F.i) Angular Fourier modes of the strain field pattern in E.ii showing peaks in the second and fourth modes.
-(F.ii) Fit of the linear elastic theory to cell strain field in the vertex model.
-The solid blue is the fitted strain field with the fit parameters $d_{0}=(10.92\pm 0.01)\times 10^{-2}a_{0}$, $\tilde{d}_{xx}=(8.29\pm 0.01)\times 10^{-2}a_{0}$, $\tilde{d}_{xy}=(2.93\pm 0.01)\times 10^{-2}a_{0}$ and $\mu=1.13\pm 0.01$.
+(F.ii) The fit of the linear elastic theory to the cell strain field in the vertex model.
+The solid blue is the fitted strain field with the fit parameters $d_{0}=0.72\pm 0.07a_{0}$, $\tilde{d}_{xx}=0.65\pm 0.07a_{0}$, $\tilde{d}_{xy}=0.00\pm 0.07a_{0}$ and $\mu=1.3\pm 0.1$.
+The uncertainty intervals for fit parameters are obtained by bootstrapping the data (see Methods).
 II.2 Viscous dissipation determines tissue relaxation dynamics
-Our analysis of the strain in the ablated tissue after it has relaxed allowed us to study its elastic properties.
-Now, we explore dynamics of tissue strain as it approaches the equilibrium, which will allow us to study dissipative processes in fly wing epithelium.
+Our analysis of the strain in the ablated tissue after it has relaxed allows us to study its elastic properties.
+Now, we explore the transient dynamics of tissue strain as it approaches the mechanical equilibrium, which will allow us to study dissipative processes in fly wing epithelium.
 We consider two possible dissipation mechanisms: viscosity and friction with a substrate.
-Interestingly, in a continuum model of a tissue as an elastic sheet, these two dissipation mechanisms lead to qualitatively different relaxation dynamics upon insertion of a force dipole.
-In particular, in an elastic sheet with only viscous dissipation, all strain field components decays as $1/r^{2}$ with the distance from the force dipole throughout the relaxation with the amplitude evolving in time, see SI and Fig.
+Interestingly, in a continuum model of tissue as an elastic sheet, these two dissipation mechanisms lead to qualitatively different relaxation dynamics upon insertion of a force dipole.
+In particular, in an elastic sheet with only viscous dissipation, all strain field components decay as $1/r^{2}$ with the distance from the force dipole throughout the relaxation with the amplitude evolving in time, see SI and Fig.
 2 (A.i).
-However, in an elastic sheet with only frictional dissipation, the strain field propagates through the sheet: at distances smaller than the propagation front, the strain field converges to the steady state solution and rapidly decays beyond the front, see SI and Fig.
+However, in an elastic sheet with only frictional dissipation, the strain field propagates through the sheet: at distances within the propagation front, the strain field converges to the steady state solution and rapidly decays beyond the front, see SI and Fig.
 2 (A.ii).
+To compare the strain field dynamics in viscous and frictional dissipation models within a cellular system, we performed vertex model simulations of laser ablation that include only the frictional dissipation and compare it with a viscous dissipation, see SI for details.
+As discussed above, the signature difference between viscous and frictional dissipation is the power law decay of the shape $1/r^{2}$ in the former and the existence of a propagation front in the latter.
+With a frictional dissipation, we indeed find a clear propagation of the strain field that converges to the steady state value as the propagation front expands (2 (B.ii) and SI Fig.
+12 (B)), consistent with the analytical calculation (2 (A.ii)).
+In contrast, with viscous dissipation, we observe that the strain field decays as $1/r^{2}$ in all time points (2 (B.i) and SI Fig.
+12 (A)).
 To quantify the dynamics of tissue relaxation following the laser ablation, we measure the strain field in time intervals of $4.5s$, starting from $1s$ after the ablation, (SI Fig.
-7).
+8).
 We find that the angular strain pattern is the same as in the mechanical equilibrium one, see Movie 1.
-We analyse the dynamics of the strain radial profile by calculating the second angular Fourier mode of the strain field at all measured timepoints, see Fig.
-2 (B.i) and SI Fig.
-7.
-We find that at all timepoints the strain field is consistent with $1/r^{2}$ decay and we do not see any signature of a propagating front, see Fig.
+We analyze the dynamics of the strain radial profile by calculating the second angular Fourier mode of the strain field at all measured timepoints, see Fig.
+2 (C.i) and SI Fig.
+8.
+We find that the strain field is consistent with $1/r^{2}$ decay and we do not see any signature of a propagating front, see Fig.
 2.
-Therefore, we conclude that the viscosity is the dominant mode of dissipation in the fly wing epithelium on timescales above $1s$, corresponding to the initial timepoint we observe after the ablation, and below $T_{e}\approx 50s$.
-To make sure that a predicted propagation front could be observed in presence of frictional dissipation we performed vertex model simulations of laser ablation that include only the frictional dissipation, see SI for details.
-We indeed find a clear propagation of the strain field that converges to the steady state value as the propagation front expands (2 (B.ii) and SI Fig.
-8), as predicted by the analystical calculation (2 (A.ii)).
-Figure 2: Time evolution the second Fourier mode $m=2$ of the strain field $xy$ component induced by a force dipole for: (A.i) linear elastic sheet with viscous dissipation, (A.ii) linear elastic sheet with frictional dissipation, (B.i) laser ablation experiment (B.ii) vertex model simulation with frictional dissipation.
-Dashed line indicates the power-law $r^{-2}$ decay for reference.
-Table 1: Parameters obtained by fitting the continuum theory to spatio-temporal pattern of strain field accumulated in the tissue after the ablation.
-exp. $d_{0}~{}[\mu m]^{2}$ $\tilde{d}_{xx}~{}[\mu m]^{2}$ $\tilde{d}_{xy}~{}[\mu m]^{2}$ $\mu$ $\tau_{1}~{}(s)$ $\tau_{2}~{}(s)$
-1 $16\pm 1$ $15\pm 1$ $2\pm 1$ $1.9\pm 0.1$ $11\pm 1$ $12\pm 1$
-2 $15\pm 2$ $31\pm 2$ $0\pm 2$ $2.1\pm 0.3$ $48\pm 2$ $55\pm 2$
+Therefore, we conclude that the viscosity is the dominant mode of dissipation in the fly wing epithelium on timescales below $T_{e}\approx 30s$.
+Figure 2: Time evolution the second Fourier mode $m=2$ of the strain field $xy$ component induced by a force dipole for: (A.i) linear elastic sheet with viscous dissipation, (A.ii) linear elastic sheet with frictional dissipation, vertex model simulation with (B.i) viscous dissipation and, (B.ii) frictional dissipation.
+(C) Laser ablation experiment.
+(C.i) Time evolution of $\tilde{U}_{xy}$ shows a radial profile as $r^{-2}$, consistent with a viscous dissipation mechanism.
+(C.ii) Quantifying the viscous relaxation of strain as described in the text, which allows the measurement of viscous relaxation timescales.
+The black dashed line indicates the power-law $r^{-2}$ decay for reference.
+Each data point represents the strain field average over $5$ smaller bins and the errors are the corresponding standard deviations over those bins.
 Our continuum model of an elastic sheet provides the dynamical strain propagator ${G}_{ij}(\bm{r},t)={G}_{ij}(\bm{r}){T}_{ij}(t)$, where no sum over repeated indices is performed.
-${G}_{ij}(\bm{r})$ is the steady state propagator, and ${T}_{ij}(t)$ is the relaxation matrix whose components contains two relaxation factors $g_{1}(t)=1-\exp{(-t/\tau_{1})}$ and $g_{2}(t)=1-\exp{(-t/\tau_{2})}$, where the two relaxation timescales
+${G}_{ij}(\bm{r})$ is the steady state propagator, and ${T}_{ij}(t)$ is the relaxation matrix whose components contain two relaxation factors $g_{1}(t)=1-\exp{(-t/\tau_{1})}$ and $g_{2}(t)=1-\exp{(-t/\tau_{2})}$, where the two relaxation timescales
 $\tau_{1}$ $=\frac{\overline{\eta}}{\overline{K}}\frac{1+2\eta/\overline{\eta}}{1+\mu}$ (3)
 $\tau_{2}$ $=\frac{\eta}{K}$ (4)
-are related to the bulk and shear elastic constants, and bulk and shear viscosities $\overline{K},2K,\overline{\eta},\text{ and }2\eta$, respectively (see SI).
-To extract $\tau_{1}$ and $\tau_{2}$ from experimental data, as described in SI and shown in SI Fig.
-1.
-We find that the two relaxation timescales are very similar, although they vary between experiments, see Table 1.
-It is at the moment unclear whether the difference between the relaxation timescales in two experiments stems from variability of elastic constants or viscosity, or some other factor.
-While this will be an important open question to resolve in future, further investigation of relaxation timescale variability is beyond the scope of this work.
-So far, our detailed analysis of the strain in the fly wing epithelium in response to the laser ablation shows that the tissue behaves as a two-dimensional viscoelastic solid.
-With this knowledge in hand, we can now infer force dipoles generated by cell divisions, and thereby study their mechanics, by simply observing the dynamics of the surrounding tissue strain caused by the division, without the need to directly perturb the tissue.
+are related to the bulk and shear elastic constants, and bulk and shear viscosities $\overline{K},2K,\overline{\eta},\text{ and }2\eta$, respectively (see SI I.B.(ii)).
+We extract $\tau_{1}$ and $\tau_{2}$ from experimental data, see Methods, SI, and SI Fig.
+13.
+We find that the two relaxation timescales are very similar: $\tau_{1}=11\pm 1s$ and $\tau_{2}=12\pm 1s$, where intervals of uncertainty correspond to the uncertainty of the fit, see Methods.
+As mentioned above, the Experiment 1 presented here was selected based on our ability to segment and track cells in a radius of $40~{}\mu m$ surrounding the laser ablation, which was not possible for most of the other experiments.
+Here, we highlight that in Experiment 2, the angular and radial pattern of the strain field are consistent with that of an elastic sheet with viscous dissipation on times comparable to the relaxation timescales $\tau_{1}$ and $\tau_{2}$ extracted above, as reported in the SI Fig.
+13.
+Interestingly, on longer times radial profile of strain departs from $r^{-2}$ and continues to evolve throughout the experiment.
+Similar additional motion is observed even in the Experiment 1 reported above, but only after about $2~{}\rm{min}$, see SI Fig.
+11.
+This suggests that additional dynamics are due to a wound-healing response in the tissue and understanding this behavior would require a dedicated experimental exploration in the future.
+In summary, our detailed analysis of the strain in the fly wing epithelium in response to the laser ablation shows that the tissue behaves as a two-dimensional viscoelastic solid.
+These results suggest that we can now infer force dipoles generated by cell divisions, and thereby study their mechanics, by simply observing the dynamics of the surrounding tissue strain caused by the division, without the need to directly perturb the tissue.
 III Elastic deformation imprinted by a cell division
 Cell division is a complex process during which the mother cell is split into two daughter cells.
-Here, we aim to characterise the mechanical forces generated by the dividing cell by analysing the strain field a dividing cell induces in the fly wing epithelium.
-We analysed cell divisions in previously published experiments Dye et al.
+Here, we aim to characterize the mechanical forces generated by the dividing cell from the analysis of the strain field that the dividing cell induces in the epithelium.
+We analyze cell divisions in previously published experiments Dye et al.
 (2021) on explanted D.
-melanogaster wing imaginal discs  [^2] Note that this developmental stage preceeds the pupal stage in which linear laser ablations were performed Piscitello-Gómez et al.
-(2023) and although both are fruit fly wing tissues, it is possible that tissue elastic constants have changed between the two stages..
+melanogaster wing imaginal discs.
+We are interested in the generic features of the response of epithelial tissues to a mechanical perturbation, such as the laser ablation or a cell division.
+Therefore, we aim to extract normalised force dipoles in the surface plane generated by cell divisions using the same approach as with the laser ablation experiments.
+However, we perform the analysis independently of any results obtained from laser ablations in the pupal wing.
+The wing disc developmental stage precedes the pupal stage in which linear laser ablations analysed above were performed Piscitello-Gómez et al.
+(2023), and their mechanical moduli may be different.
+However, a quantification of the strain field generated by cell divisions is more convenient in the wing disc stage compared to the pupal wing stage.
+Cell divisions in the pupal wing differ from divisions in the wing disc in that they are patterned in time and space, with most cells dividing at least once within several hours and the tissue exhibits strong tissue shear flows and active cell rearrangements Etournay et al.
+(2016).
+Therefore, analysis of the strain field induced by individual divisions is practically impossible in the pupal wing due to additional dynamics in the tissue.
+However, in both tissues, the apical surfaces are mechanically special as they contain a high concentration of actomyosin cortex and adherens junctions between cells that are important for morphogenesis.
+Therefore, it is likely that apical mechanics plays a significant role in the overall tissue mechanics in both tissues.
 Wing discs were cultured over about $13$ hours, during which the tissue was imaged every $5$ minutes, and individual cells were segmented and tracked over time.
@@
 Before the actual cell division, the mother appears to inflate Fig.
-3 (A.ii), which is not a real increase in cell size, but is rather a consequence of the cell nucleus arriving to the apical surface of the cell to divide Gibson et al.
+3 (A.ii).
+This is not a real increase in cell size, but is rather a consequence of the cell nucleus arriving at the apical surface of the cell to divide and transiently stretching the cell apical surface Gibson et al.
 (2011); Strzyz et al.
@@
 Fig.
-3 (A.iv) qualitatively summarizes these experimental observations.
+3 (A.iv) qualitatively summarizes the experimental observations.
 For each cell division, we quantify the tissue strain field in the vicinity of the dividing cell, as described in Section II, relative to the starting timepoint $t_{0}=-60~{}\rm{min}$ before the cell division time.
@@
 3 (B).
-The timescale on which the cell division forces evolve, which is on the order of tens of minutes, is much larger than the relaxation timescale extracted in Section II, which is below one minute.
-Therefore, the observed strain field dynamics is largely due to changes of the force dipole generated by the cell division, and not due to much faster stress propagation through the tissue.
+The timescale on which the cell division forces evolve, which is on the order of tens of minutes, is much larger than the minutes relaxation timescale reported in the wing disc Dye et al.
+(2021).
+Therefore, the strain field dynamics observed over tens of minutes is largely due to changes of the force dipole generated by the cell division, and not due to much faster stress propagation through the tissue.
+Although it is possible that on these longer timescales cells adapt to the experienced stress and adapt their reference shape, a clear positive correlation between stress and strain has been reported in our past work Dye et al.
+(2021).
+Therefore, it is reasonable to interpret the change in cell elongation as a measure of elastic strain, even if the adaptation to stress effectively renormalizes the elastic constants, as discussed in Ref.
+Dye et al.
+(2021).
 Figure 3: (A) Snapshots of a cell division in three stages.
-(A.i) As the mother cell starts the division process, (A.ii) it rounds up, corresponding to the inflation, then (A.iii) it shrinks and divides into the daughter cells.
+(A.i) As the mother cell starts the division process, (A.ii) it rounds up, corresponding to the inflation, and then (A.iii) it shrinks and divides into the daughter cells.
 The red dots show the mother and the daughter cells.
@@
 (C) Ensemble average ($N=682$) of $\tilde{U}_{xy}$ around the dividing cell: (C.i) at the peak of the inflation $t=-20~{}[\rm{min}]$ (corresponding to (A.ii) and (B.ii)), and (C.ii) after the division $t=30~{}[\rm{min}]$ (corresponding to (A.iii) and (B.iii)).
-(D) The force dipole generated by a cell division as a function of time.
-(E) A schematic summary of the force dipoles during stages of a cell division.
+The radial extent of the plots is $6\mu m$.
+(D) The angular Fourier transform of the $\tilde{U}_{xy}$ at a distance of $r=3~{}\mu m$ from the center of the mother cell at $t=-20~{}[\rm{min}]$ (D.i) and at $t=30~{}[\rm{min}]$ (D.ii).
+Figure 4: (A) The fit of linear elastic theory to the measured strain field around cell divisions at (A.i) the peak of inflation $t=-20~{}[\rm{min}]$ and (A.ii) after division $t=30~{}[\rm{min}]$.
+Each data point shows the average strain field in polar bins and the error bars are the standard deviation over these bins.
+(B) The force dipole generated by a cell division as a function of time.
+(C) A schematic summary of the force dipoles during stages of a cell division.
 We find that a single cell division generates weaker forces compared to those induced by laser ablation (compare Fig.
@@
 Therefore, we average the strain field from multiple cell divisions ($N=682$).
-We first align the centers of the cell division, defined as the area weighted average of daughter cell centers at the cell division time ($t=0$).
+We first align the centers of the cell division, defined as the area-weighted average of daughter cell centers at the cell division time ($t=0$).
 We next rotate each division such that the orientation of the line connecting daughter cell centers is oriented along the $x-$axis at the cell division time.
 We show the resulting strain field component $\tilde{U}_{xy}$ in Fig.
-3 (C) at times $t=-20~{}\rm{min}$, corresponding to the peak of the cell inflation, and $t=30~{}\rm{min}$ after the cell division.
+3 (C) and the corresponding angular Fourier modes (D) at times $t=-20~{}\rm{min}$, corresponding to the peak of the cell inflation, and $t=30~{}\rm{min}$ after the cell division.
 In Fig.
@@
 In Fig.
-3 (C.ii), the pattern is closer to 4-fold symmetric one, such as would be induced by an anisotropic force dipole.
+3 (C.ii), the pattern is closer to a 4-fold symmetric one, such as would be induced by an anisotropic force dipole.
+Finally, we find that radial profiles of the Fourier modes are consistent with $r^{-2}$ decay (SI Fig.
+15), as was the case with laser ablation experiments above.
 We fit the linear elastic theory Eq.
 2 to the experimentally measured strain field at each time point and extract the resulting force dipole $d_{ij}(t)$.
+Comparison between the fit and the data is shown for the $xy$ component of the strain field at $t=-20\rm{min}$ and $t=30\rm{min}$ in Fig.
+4 B.
 We find that the isotropic force dipole component $d_{0}$ rises sharply during the mother cell inflation and peaks at about $t=-20~{}\rm{min}$ (magenta line in Fig.
-3 D).
+4 B).
 However, by the time of the new bond formation, the isotropic component has disappeared and remains negligible at later times.
-This result shows that cell does not noticeably grow during the division, but rather that the division reorganises the available material into the two daughter cells.
+This result shows that cells do not noticeably grow during the division, but rather that the division reorganizes the available material into the two daughter cells.
 The traceless part of the force dipole $\tilde{d}_{xx}$ grows as the new bond is formed and persists in the tissue multiple hours after the division (green line in Fig.
-3 D), indicating that the division-induced strains remain imprinted in the tissue long after the division has been executed.
+4 B), indicating that the division-induced strains remain imprinted in the tissue long after the division has been executed.
 Over time, $\tilde{d}_{xx}$ slowly decays while its fluctuations increase, representing the effect of other cell divisions and rearrangements that appear in the tissue.
 We estimate the tissue fluidization time from the decay of $\tilde{d}_{xx}$ in time by fitting an exponential function to it.
-We find the fluidization timescale to be $\tau_{\alpha}\simeq 3.5\pm 0.7$ minutes, over which $\tilde{d}_{xx}$ vanishes and, therefore, the strain field induced by the division is erased.
+We find the fluidization timescale to be $\tau_{\alpha}\simeq 3.5\pm 0.7$ hours (see Methods), over which $\tilde{d}_{xx}$ vanishes and, therefore, the strain field induced by the division is erased.
 Note that component $\tilde{d}_{xy}$ remains negligible at all times.
@@
 Since the mother cell attains an almost perfectly circular shape during the inflation, any initial anisotropy of the mother cell at time $t_{0}$ gives rise to an anisotropic force dipole component during the inflation.
-Therefore, the observed negative value of $\tilde{d}_{xx}$ indicates that mother cells are slightly elongated along the future division axis, consistent with the well known Hertwig’s rule Hertwig (1884); Théry et al.
+Therefore, the observed negative value of $\tilde{d}_{xx}$ indicates that mother cells are elongated along the future division axis, consistent with the well-known Hertwig’s rule Hertwig (1884); Théry et al.
 (2007); Gibson et al.
@@
 We summarise these results in Fig.
-3 (E).
+4 (C).
 IV Discussion
 In this work, we devised a full spatio-temporal analysis of tissue strain induced by two types of force dipole fields: linear laser ablations and cell divisions.
-In this way, we have shown that the developing Drosophila wing epithelium behaves as a two-dimensional elastic solid with viscous dissipation.
+In this way, we have shown that the developing Drosophila wing epithelium can be described as a two-dimensional elastic solid with viscous dissipation.
 We also measured its two-dimensional Poisson ratio, which we find to be negative, consistent with the values reported in the literature Dye et al.
@@
 (2023).
+Our analysis was based on laser ablation of the wing tissue during pupal development, at which time the wing consists of two apposed epithelial layers.
+Although the experimental observations are consistent with a model of the wing apical surface as a linear elastic sheet, in the future it might be interesting to probe the three-dimensional structure response of the tissue with deeper ablations and in wings where the two epithelial layers have been separated Sun et al.
+(2021).
 We characterized the dynamics of mechanical forces generated by a dividing cell throughout the cell division process.
-We found that strain fields imprinted by a cell division are slowly erased on the timescale of $\tau_{\alpha}=3.5\pm 0.7~{}\rm{hr}$, which can be understood as a fluidization timescale of the Drosophila wing tissue.
-Furthermore, our analysis shows that there is no cell growth during a division and stresses induced by the cell division correspond to those of a pure shear force dipole.
+We found that the strain fields imprinted by a cell division are slowly erased on the timescale of $\tau_{\alpha}=3.5\pm 0.7~{}\rm{hr}$, which can be understood as a fluidization timescale of the Drosophila wing tissue.
+Furthermore, our analysis shows that there is no cell growth during a division, and stresses induced by the cell division correspond to those of a pure shear force dipole.
 Finally, we found that Hertwig’s rule applies in the D.
@@
 An important implication of our work is that the Drosophila wing epithelium is a solid whose fluidization time $\tau_{\alpha}$ is not negligible on the developmental timescales.
-For example, large scale shape changes that the wing undergoes during eversion are executed on a similar timescale Fuhrmann et al.
+For example, large-scale shape changes that the wing undergoes during eversion are executed on a similar timescale Fuhrmann et al.
 (2023).
-A possible physical picture of the Drosophila wing epithelium that could account for our observations, is that of a glassy material, or as was put forward in glass literature, a solid that flows Dyre (2023); Ozawa and Biroli (2023); Tahaei et al.
+A possible physical picture of the Drosophila wing epithelium that could account for our observations is that of a glassy material, or as was put forward in the glass literature, a solid that flows Dyre (2023); Ozawa and Biroli (2023); Tahaei et al.
 (2023).
@@
 In summary, our results quantify force dipoles generated by dividing cells in developing D.
-melanogaster wing epithelia, as well es elastic interactions between cells, which paves the way to development tissue mechanics theory and motivates further experimental study of collective cell dynamics.
+melanogaster wing epithelia, as well as elastic interactions between cells, which paves the way to the development of tissue mechanics theory and motivates further experimental study of collective cell dynamics.
 Acknowledgements
-We would like to thank Michael Staddon, Frank Jülicher, Jonas Neipel for usefull discussions and comments.
-NAD, JFF, and GC acknowledge funding from the Max Planck Society, Excellence Cluster Physics of Life (EXC-2068-390729961), Deutsche Forschungsgemeinschaft (SPP1782/EA4/10-2), Deutsche Krebshilfe/MSNZ Dresden.
+We would like to thank Michael Staddon, Frank Jülicher, Jonas Neipel for useful discussions and comments.
+Funding for the biological experiments was supported by the Max Planck Society and the Deutsche Forschungsgemeinschaft (Cluster of Excellence Physics of Life, EXC-2068-390729961 and SPP1782/EA4/10-2).
+NAD additionally acknowledges funding from Deutsche Krebshilfe/MSNZ Dresden.
+MP and ND acknowledge funding by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) - Project number DY 180/2-1 (eBer-24-23102) and PO 3023/2-1 - 544201605.
 V Methods
 V.1 Cell segmentation and tracking
-The linear ablation experiments came from the pupal morphogenesis stage and were previously published in Ref.
+The linear ablation experiments were performed during the pupal morphogenesis stage and were previously published in Ref.
 Piscitello-Gómez et al.
-(2023), where the authors analyzed initial retraction velocity of the ablated bonds.
-Here, we took these experiments and fully segmented and tracked cells in the entire field using the FIJI plugin, TissueAnalyzer Aigouy et al.
+(2023), where the authors analyzed the initial retraction velocity of the ablated bonds.
+Here, we re-analyzed movies of those experiments, and fully segmented and tracked cells in the entire field using the FIJI plugin, TissueAnalyzer Aigouy et al.
 (2016), using a bond cutoff of 2 pixels.
 Then, after the manual correction of segmentation errors, we use TissueMiner Etournay et al.
-(2016) to analyse changes in the cell shape and size.
-The cell division experimental data came from the wing disc stage and was previously segmented, tracked and analyzed in Ref. Dye et al. (2017, 2021).
+(2016) to analyze changes in the cell shape and size.
+The cell division experimental data came from the wing disc stage and was previously segmented, tracked, and analyzed in Ref. Dye et al. (2017, 2021).
 V.2 Tissue relaxation after a laser ablation
-After the ablation, nearby tissue relaxes to the new mechanical equilibrium.
-However, beyond $2$ minutes after the ablation we observed additional motion of the tissue which we interpret as the onset of the would healing response in the tissue.
-In this study, we are interested in the stress released by the ablation in the tissue, not the wound healing process, and therefore we limit our analysis to times up to $2$ minutes after the laser ablation.
+After the ablation, nearby tissue relaxes to a new mechanical equilibrium.
+In Experiment 1 reported in the main text, we estimate equilibrium to be reached after about $30s$ after ablation, when the tissue does not visibly evolve any more.
+However, this equilibrium state is not permanent, as at about $1$ minute after the ablation, we observe additional motion of the tissue, which we interpret as the onset of the wound-healing response.
+In this work, we are interested in the stress response to the ablation, not the wound-healing process, and therefore we limit our analysis to times below $1$ minute after the laser ablation.
+Note that in Experiments 2-4, reported in detail in SI, the additional patterns appear to emerge earlier and we could not extract as clear an intermediate regime as in Experiment 1.
+However, even in these experiments the profiles of strain observed during the relaxation are consistent with the linear elastic theory, thereby further strengthening our results from Experiment 1.
 V.3 Tissue strain field measurement
 Strain field consists of isotropic strain $U_{kk}(\mathbf{r},t)/2$ and shear strain $\tilde{U}_{ij}(\mathbf{r},t)$.
-To measure the strain field in the tissue we calculate the change in cell shape and size as described in Merkel et al.
+To determine the strain field from the changes of cell shape and area, we use the method based on the cell triangulation Etournay et al.
+(2015); Merkel et al.
 (2017); Duclut et al.
-(2021) at each timepoint.
-Since there are no cell rearrangements and divisions we estimate the accumulated tissue shear strain field in a cell from the difference between the cell shape tensors
+(2021) in which the co-rotational change of the cell elongation tensor $Q_{ij}$ corresponds to the tissue shear rate, and changes in the logarithm of cell area to the isotropic strain rate.
+Here, we approximate the accumulated tissue strain by the change in cell elongation tensor and in the logarithm of cell area for pure shear and isotropic components, respectively.
+To define the cell elongation tensor $Q_{ij}$, we choose the sub-cellular triangulation of the tissue obtained by connecting two adjacent vertices to the geometric center of cells, as described in Duclut et al.
+(2021).
+We estimate the accumulated tissue shear strain field in a cell from the difference between the cell shape tensors
 $\tilde{U}_{ij}(t)$ $\simeq Q_{ij}(t)-Q_{ij}(0)$ (5)
-The only approximation is that we neglect to account for change in cell elongation due to local rotations.
-We determined the exact cell shear tensor at several timepoints and we found that the difference from our approximation are small.
+The only approximation is that we neglect to account for changes in cell elongation due to local rotations.
+We determined the exact cell shear tensor at several timepoints, and we found that the differences from our approximation are small.
 Therefore, we use the approximate expression for simplicity.
-We estimate the isotropic strain field in a cell from the change in cell area $a$
-$U_{kk}(t)$ $=\log{\left(\frac{a(t)}{a(0)}\right)}\quad.$ (6)
-We define the tissue strain field $U_{ij}(\mathbf{r},t)$ at position $\mathbf{r}$ to have the value of the strain tensor obtained for the cell inside of which point $\mathbf{r}$ lies.
+The isotropic strain field in a cell is defined as the accumulated relative area change rate Etournay et al.
+(2015); Merkel et al.
+(2017), which we measure relative to the initial cell state as
+$U_{kk}(t)$ $=\log{\left(\frac{a(t)}{a(0)}\right)}\quad,$ (6)
+where $a$ is the cell area.
+Finally, we define the tissue strain field $U_{ij}(\mathbf{r},t)$ at position $\mathbf{r}$ to have the value of the strain tensor obtained for the cell inside of which point $\mathbf{r}$ lies.
 V.4 Discrete angular Fourier transform
-To investigate the angular pattern of the strain field and study its radial profiles, we look at the angular discrete Fourier transformation of the tissue strain field.
-The strain in position $\bm{r}$ is given by ${U}_{ij}(r,\varphi)$ where $r$ is the distance to the center of the cut and $\varphi$ is the polar angle.
+We examine the angular discrete Fourier transformation of the tissue strain field to investigate its angular pattern and study its radial profiles.
+The strain in position $\bm{r}$ is represented by ${U}_{ij}(r,\varphi)$, where $r$ is the distance to the center of the cut and $\varphi$ is the polar angle.
 We use the notation ${U}_{ij}(r,m)$ for the $m$-th angular mode of discrete angular Fourier transform at a distance $r$, and it is defined as,
 ${U}_{ij}(r,m)=\frac{1}{N}\sum_{i=1}^{N}{U}_{ij}(r,\varphi_{i})e^{im\varphi_{i}},$ (7)
-where $\varphi_{i}$s are $N$ uniformly distributed number between 00 and $2\pi$.
+where $\varphi_{i}$ are $N$ uniformly distributed numbers between 00 and $2\pi$.
+In rare cases a point with coordinates $(r,\varphi_{i})$ does not belong to a cell, making ${U}_{ij}(r,\varphi)$ undefined at that point.
+To handle such cases, we modify the Fourier transform as follows,
+${U}_{ij}(r,m)=\frac{\sum_{i=1}^{N}{U}_{ij}(r,\varphi_{i})e^{im\varphi_{i}}\iota(r,\varphi_{i})}{\sum_{i=1}^{N}\iota(r,\varphi_{i})},$ (8)
+where $\iota(r,\varphi_{i})=1$ if the point belongs to a cell and 00 otherwise.
+[…]
 V.5 Fit of the elastic theory to data
-We measure force dipole, the ratio of shear modules to bulk modules, and viscosity by fitting the strain field pattern to the elastic and viscoelastic theories. First, we define a cost function
-$E=\sum_{\alpha\in\rm{cells}}\Big{[}(U_{kk}^{\alpha}-u_{kk}(\bm{r}^{\alpha})^{2}+(\tilde{U}_{xx}^{\alpha}-\tilde{u}_{xx}(\bm{r}^{\alpha})^{2}+\\ (\tilde{U}_{xy}^{\alpha}-\tilde{u}_{xy}(\bm{r}^{\alpha})^{2}\Big{]},$ (8)
-where ${U}_{ij}^{\alpha}$ is the cell strain measured from data, $\bm{r}^{\alpha}$ is the position of centre of each cell and ${u}_{ij}(\bm{r}^{\alpha})$ is the strain field predicted from theory in position $\bm{r}^{\alpha}$.
-By minimising $E$, we find the force magnitude of the force dipole, the ratio of the shear modulus to bulk modulus, and the relaxation timescales.
-In the elastic case, first, we minimise a steady state signal to find $\mu$ then we fix $\mu$ and repeat the minimisation process to find ${d}_{ij}$.
-In viscoelastic case, we use $\mu$ and ${d}_{ij}$ from the steady state elastic response, then we minimise $E$ at different time points to find $T_{ij}(t)$.
-To analyse cell divisions, we first choose timepoint $t=25~{}\rm{min}$ and change the reference time $t_{0}$ from $-120$ to $-40$.
-We fit the data for different $t_{0}$ in that range and obtain corresponding set of $\mu(t_{0})$ values.
-We determine the parameter $\mu$ as the average $\langle\mu(t_{0})\rangle$ and we do not vary it in subsequent fitting of the strain field.
-Then, to determine the force dipole at each timepoint $t$ we fix $t_{0}=-60~{}\rm{min}$ fit the theory to the data to find the force dipole at different time points.
+We measure the force dipole, the ratio of shear moduli to bulk moduli, and viscosity by fitting the strain field pattern to the elastic and viscoelastic theories. First, we define a cost function
+$E=\sum_{\alpha\in\rm{cells}}\Big{[}(U_{kk}^{\alpha}-u_{kk}(\bm{r}^{\alpha}))^{2}+(\tilde{U}_{xx}^{\alpha}-\tilde{u}_{xx}(\bm{r}^{\alpha}))^{2}+\\ (\tilde{U}_{xy}^{\alpha}-\tilde{u}_{xy}(\bm{r}^{\alpha}))^{2}\Big{]},$ (9)
+where ${U}_{ij}^{\alpha}$ is the cell strain measured from data, $\bm{r}^{\alpha}$ is the position of the center of each cell and ${u}_{ij}(\bm{r}^{\alpha})$ is the strain field predicted from theory at position $\bm{r}^{\alpha}$.
+By minimizing $E$, we find the magnitude of the force dipole, the ratio of the shear modulus to bulk modulus, and when fitting the relaxation dynamics, the relaxation timescales.
+When fitting the steady state strain field measured in Experiment 1, we wanted to take into account the fluctuations in the steady state.
+For this, we sampled the strain field at 5 timepoints uniformly spaced between $36s$ and $58.5s$ after ablation.
+We constructed datasets by bootstrapping the sampled data $10$ times, and we fit the linear elastic theory for each dataset.
+The reported parameter values and uncertainties are mean values and standard deviations of these 10 fit outcomes.
+To fit the relaxation dynamics and extract the relaxation timescale, we consider the strain field at each time point $t_{i}$, spaced $4.5s$ apart, during the relaxation and use the steady state parameters and force dipole to fit the elements of the matrix $\bm{T}(t_{i})$.
+Finally, we fit an exponential function to the obtained matrix element values at different time points, Fig.
+2.C.ii, and report the relaxation timescale values and their uncertainties obtained by this exponential fit.
+To analyze cell divisions, we first reduce noise in the strain field by applying a moving average over a $10$-minute window.
+Next, we select the time point $t=30\rm{min}$ and systematically vary the reference time $t_{0}$ from $-120\rm{min}$ to $-40\rm{min}$.
+For each $t_{0}$ within this range, we fit the steady state linear elastic theory to the measured strain field to extract the corresponding set of $\mu(t_{0})$ values.
+The parameter $\mu$ is then determined as the average $\langle\mu(t_{0})\rangle$, which remains fixed for the subsequent strain field analyses (see SI Fig.
+14).
+This is done to more precisely measure the parameter $\mu$, which cannot be extracted independently of $d_{0}$ when traceless symmetric components of the force dipole vanish.
+Namely, in that case it is impossible to factorize the product $\mu D_{0}$ as the two factors only appear in the form of the product, see SI Eqs.
+9 and 10.
+Finally, to compute the force dipole at each timepoint $t$ we set $t_{0}=-60\rm{min}$, and we perform the fit at each time point using the bootstrapping algorithm over the ensemble of all divisions, obtaining the force dipole values and their uncertainties at each time point, while using the previously obtained values of $\mu$ as a constant.
 V.6 Estimation of the tissue fluidization time
-To measure the decay time scale of the anisotropic force dipole $\tilde{d}_{xx}$, we fit an exponential function to normalized force dipole $y=\tilde{d}_{xx}(t)/\langle a\rangle$ as $f(y)=\tilde{d}_{xx}(t=0)\exp[-t/\tau_{\alpha}]$ to observed data in Fig.
-3 (D).
-From that, we obtain the fluidization times scale $\tau_{\alpha}=3.5\pm 0.7\rm{hr}$ and $\tilde{d}_{xx}(t=0)/\langle a\rangle=0.041\pm 0.001$.
+To measure the decay timescale of the anisotropic force dipole $\tilde{d}_{xx}$, we fit an exponential function to normalized force dipole $y=\tilde{d}_{xx}(t)/\langle a\rangle$ as $f(y)=\tilde{d}_{xx}(t=0)\exp[-t/\tau_{\alpha}]$ to observed data in SI Fig.
+16.
+From that, we obtain the fluidization timescale $\tau_{\alpha}=3.5\pm 0.7\rm{hr}$ and $\tilde{d}_{xx}(t=0)/\langle a\rangle=0.53\pm 0.01$.
+The reported errors correspond to the range of parameters over which an exponential function successfully fits the data.